# Promptly Predicting Structures: The Return of Inference

### v1

@@ 2 Related Work
Predicting structures with prompts is challenging; the output correspond to multiple mutually-constraining decisions. Structures that can be flattened into label sequences can be cast as text generation tasks amenable to the LLM interface. Blevins et al. (2023) present a structured prompting paradigm to generate labels in an auto-regressive fashion for part-of-speech tagging, named entity recognition, and sentence chunking tasks. They show that this approach works well in a few-shot setting. Such tasks are also amenable for templated slot filling in a few shot setting Cui et al. (2021). Liu et al. (2022) propose an autoregressive structured prediction framework that reformulates the problem into predicting a series of structure-building action steps.
Structured prediction tasks can be made prompt-friendly by identifying output components and framing each one into natural instructions He et al. (2015); Du and Cardie (2020) using the QA format. Mekala et al. (2023) decompose semantic parsing into a series of questions. However, they do not use inference and instead, aggregate outputs sequentially. Yang et al. (2022) use QA prompts to classify pairs of entity mentions as coreferent or not without aiming to form entity clusters. The near contemporary work of Lin et al. (2023) presents a prompt-driven heuristic for event type classification. Wang et al. (2023) use a programming language-based prompt and generate approach with GPT-3 for zero and few-shot event argument extraction. Recently, Le and Ritter (2023) show two techniques based on QA and document infilling for zero-shot coreference resolution that rely on recent LLMs’ ability to ingest entire documents as input.
@@ 4.1 Semantic Role Labeling
We consider two QA-SRL datasets for our experiments. 1) QA-SRLHe et al. (2015)contains 10.8k QA pairs for 4440 predicates in 1959 sentences across splits.333He et al. (2015) provide datasets in newswire and Wikipedia domains. We use the latter as it is larger. 2) FitzGerald et al. (2018)present a much larger crowd-sourced dataset—QA-SRL 2.0—containing 265k QA pairs from 64k sentences for 133k verbal predicates annotated using the QA-SRL annotation schema.
@@
Given the graph, we can perform a shortest-path search from the leftmost to the rightmost nodes; the edges in the shortest path gives us non-overlapping labels. However, doing so will generate null sequences that have a zero score. We avoid this issue by extending the algorithm to return the top shortest paths using Yen’s -shortest path algorithm Yen (1971). Subsequently, the highest scoring path among the paths which assigns exactly one span per semantic role is the optimal structure that satisfies all task constraints. For the example in Figure 2, our inference algorithm would return ‘Elrond’, ‘Aragorn’, and ‘the sword’ as the arguments for roles a, b and c respectively.666See Appendix F for a worked-out example.
@@
We evaluate at both the question-level and the structure-level. At each level, we use exact and head match metrics (Palmer et al., 2010, p. 49). At the question-level, predicted spans are compared with gold spans. For the Exact accuracy metric, answers are considered correct if the spans exactly match. Similarly, answers are considered correct if the phrasal heads of the spans match to compute the Head accuracy metric. A structure is considered accurate if predicted spans for all constituent roles match. As at the question-level, a match can be determined in an exact (Exact), or head (Head) fashion. Additionally, we also evaluate inconsistency percent () by measuring how often constraints are violated. We compare every pair of outputs(i.e., argument spans) in a predicted structure. A violation occurs when two output spans overlap.
@@ 4.2 Coreference Resolution
We use two metrics: i) F1that measures performance of the binary decisions at the question level ii) CoNLL scorePradhan et al. (2014)that averages three cluster evaluation metrics: MUC , B, and CEAF. Additionally, we measure the inconsistency percent (): the number of times constraint (3) is violated divided by the number of times the antecedent is true. This is the same as the ‘conditional violation’ defined by Li et al. (2019), and measures the fraction of times a third link between three mentions does not exist, when the other two links between them exist.
@@ 5.1 Semantic Role Labeling
Tables 1 and 2 show the results for QA-SRL and QA-SRL 2.0 respectively. We see that the constrained models produce near-consistent gains over their unconstrained counterparts. Flan-T5 does not outperform the T5 model, but iterative prompting with inference significantly improves its performance. More importantly, without inference, the models produce highly inconsistent predictions with a lowest inconsistency percent of 34%! In addition, we also provide results on GPT-4 with the constraints verbalized as system instruction.111111GPT-4 results are meant to contextualize our results with a current SOTA system. The datasets we considered may have been utilized to train GPT-4. GPT-4 also has multiple orders of magnitude more parameters than the models we consider. For GPT-4, all the questions are provided at once. Refer Appendix D.1.2 for details.
@@
| System | Head() | Head() | () |
@@
| System | Head() | Head() | () |
@@ 6.1 Model Size
We study the performance and consistency of models over varying sizes. For SRL, we compare the small, base, large, 3 billion, and 11 billion T5 models. Figure 4 shows the Head metric over T5 model sizes for QA-SRL 2.0, along with the inconsistency percent. For coreference, we compare the large, 3 billion and 11 billion variants of the Macaw model. We consider the All-Link inference for the constrained models. We report the F1 score and the inconsistency percent on the OntoNotes dataset in Figure 5. The inconsistency is shown only for the unconstrained models; the constrained ones are guaranteed to be consistent.
Across the datasets, we see that performance almost always increases and constraints provide steady gains. In QA-SRL 2.0, inconsistency reduces with increasing model size. The trend reverses for OntoNotes. Interestingly, with QA-SRL 2.0, a constrained model of a certain size often yields comparable, if not better, results as an unconstrained model of the immediately bigger model, showing that constraints might give performance gains equivalent to using a bigger model. Appendix C shows model size results for the other datasets.
@@ 6.2 Few-shot Prompting
So far, we have examined our framework in the zero-shot setting. Do the conclusions hold for few-shot settings as well? To answer this question, we use the Flan-T5-XL model which is known to yield promising results with even a few in-context examples.151515An interesting question: What counts as a shot? All questions in a structure or just one? We consider the latter. We conduct few-shot experiments for the SRL task on the QA-SRL (Figure 6) and QA-SRL 2.0 (Figure 7). We observe a clear increase in performance with increasing shots for QA-SRL. Recall that Flan T5, in a zero-shot setting for this dataset, underperforms compared to T5. Hence, the gains are not surprising: only the 5-shot constrained system beats the zero-shot constrained T5 model (38.75 Head). The model also becomes more consistent with more examples. For QA-SRL 2.0, the performance and the consistency largely remain invariant with number of shots. Constrained systems produce more accurate structures than their unconstrained counterparts.
@@ Appendix A Detailed SRL Results
| System | Exact() | Exact() | Head() | Head() | () |
@@
| System | Exact() | Exact() | Head() | Head() | () |
@@ B.1 Zero Shot Coreference Results
The complete zeros shots results for Macaw-3B and Flan T5-XL models are given in Table 8. One might notice that the Flan model is highly inconsistent as compared to the Macaw model. Note that in each case, the divisor is different since it is the number of times the antecedent is activated in equation 3. Examining the predictions, we saw that the value of the divisor for the Macaw models is roughly four times that of the Flan models. This implies that Flan tends to predict ‘No’ more often.
| Dataset | System | F1() | MUC() | B() | CEAF() | CoNLL() | () |
@@ B.2 Context Styles for Coreference Resolution
As mentioned in §5.2, for any given question pertaining to a mention-pair, we consider only the sentences where the constituent mentions occur. One can argue that surrounding context can play an important role in deciding whether or not two mentions are co-referent. Furthermore, intermediate context between the sentences can be pivotal in establishing otherwise loosely connected sentences appearing a few sentences apart. This begs the question about how much, if any, does providing the entire document help.161616For Ontonotes and GENIA, the document size went beyond the input size which is why we do not test along this dimension. On a different note, mentions can be highlighted with asterisks(as used for WSC in Raffel et al. (2020)) to explicitly state where in the context they occur. This approach can help deal with pronominal mentions which are bound to repeat. We present results when the entire document (Full) acts as context, when mentions are highlighted (Hlght), and a combination of both (Full + Hlght) in comparison to the context style where only the relevant sentences are considered (Rel). We provide detailed results for performance and consistency over differing context styles in Table 9. We see that results are comparable across all context styles.
| Dataset | System | F1() | MUC() | B() | CEAF() | CoNLL() | () |
@@ Appendix C Additional Model Size Experiments
The analysis over model sizes for QA-SRL is given in Figure 8. We see similar analysis as observed with QA-SRL 2.0 dataset. Performance improves with increasing model size while inconsistency reduces. Furthermore, constraints provide steady gains over their unconstrained counterparts.
@@ D.1.1 Iterative Prompting Template
In the iterative prompting template, we prompt the model one question at a time such that the answer for every preceding prompt is included in subsequent prompts. Assume a predicate in a sentence to have questions: {q,q,…,q}. In the first step, we prompt the model for q using the prompt template mentioned in the previous sub-section. Next, we follow with the prompt the model for q such that the question and the generated answer for q are included in the prompt. Consequently, we use a different prompt template that contains this pair and an instruction that asks the answers to not overlap. Subsequently, the prompt for q will contain the question-answer pair q and q, and so on. An example is shown in Figure 11.
@@ Appendix F SRL Inference Worked-out Example
| Rank | Role | ||
| a | b | c | |
| 1 | “Elrond” (2) | “Aragorn” (5) | “Aragorn the sword” (5) |
| 2 | “Elrond gave” (1) | “Elrond” (3) | “the sword” (4) |

### v2

@@ 2 Related Work
Predicting structures with prompts is challenging; the output corresponds to multiple mutually-constraining decisions. Structures that can be flattened into label sequences can be cast as text generation tasks amenable to the LLM interface. Blevins et al. (2023) present a structured prompting paradigm to generate labels in an auto-regressive fashion for part-of-speech tagging, named entity recognition, and sentence chunking tasks. They show that this approach works well in a few-shot setting. Such tasks are also amenable for templated slot filling in a few shot setting Cui et al. (2021). Liu et al. (2022) propose an autoregressive structured prediction framework that reformulates the problem into predicting a series of structure-building action steps.
Structured prediction tasks can be made prompt-friendly by identifying output components and framing each one into natural instructions He et al. (2015); Du and Cardie (2020) using the QA format. Mekala et al. (2023) decompose semantic parsing into a series of questions. However, they do not use inference and instead, aggregate outputs sequentially. Yang et al. (2022) use QA prompts to classify pairs of entity mentions as coreferent or not without aiming to form entity clusters. The near contemporary work of Lin et al. (2023) presents a prompt-driven heuristic for event type classification. Wang et al. (2023) use a programming language-based prompt and generate approach with GPT-3 for zero and few-shot event argument extraction. Recently, Le and Ritter (2023) show two techniques based on QA and document infilling for zero-shot coreference resolution that rely on recent LLMs’ ability to ingest entire documents as input. The concurrent work of Rajaby Faghihi and Kordjamshidi (2024) introduces an inference-time method for enforcing structural consistency that incorporates local decisions from different models using various normalizer functions.
@@ 4.1 Semantic Role Labeling
We consider two QA-SRL datasets for our experiments. 1) QA-SRLwikiHe et al. (2015)contains 10.8k QA pairs for 4440 predicates in 1959 sentences across splits.333He et al. (2015) provide datasets in newswire and Wikipedia domains. We use the latter as it is larger. 2) FitzGerald et al. (2018)present a much larger crowd-sourced dataset—QA-SRL 2.0—containing 265k QA pairs from 64k sentences for 133k verbal predicates annotated using the QA-SRL annotation schema.
@@
Given the graph, we can perform a shortest-path search from the leftmost to the rightmost nodes; the edges in the shortest path give us non-overlapping labels. However, doing so will generate null sequences that have a zero score. We avoid this issue by extending the algorithm to return the top shortest paths using Yen’s -shortest path algorithm Yen (1971). Subsequently, the highest scoring path among the paths which assigns exactly one span per semantic role is the optimal structure that satisfies all task constraints. For the example in Figure 2, our inference algorithm would return ‘Elrond’, ‘Aragorn’, and ‘the sword’ as the arguments for roles a, b and c respectively.666See Appendix F for a worked-out example.
@@
We evaluate at both the question-level and the structure-level. At each level, we use exact and head match metrics (Palmer et al., 2010, p. 49). At the question-level, predicted spans are compared with gold spans. For the Exactq accuracy metric, answers are considered correct if the spans exactly match. Similarly, answers are considered correct if the phrasal heads of the spans match to compute the Headq accuracy metric. A structure is considered accurate if predicted spans for all constituent roles match. As at the question-level, a match can be determined in an exact (Exacts), or head (Heads) fashion. Additionally, we also evaluate inconsistency percent () by measuring how often constraints are violated. We compare every pair of outputs(i.e., argument spans) in a predicted structure. A violation occurs when two output spans overlap.
@@ 4.2 Coreference Resolution
We use two metrics: i) F1that measures performance of the binary decisions at the question level ii) CoNLL scorePradhan et al. (2014)that averages three cluster evaluation metrics: MUC , B3, and CEAFe. Additionally, we measure the inconsistency percent (): the number of times constraint (3) is violated divided by the number of times the antecedent is true. This is the same as the ‘conditional violation’ defined by Li et al. (2019), and measures the fraction of times a third link between three mentions does not exist, when the other two links between them exist.
@@ 5.1 Semantic Role Labeling
Tables 1 and 2 show the results for QA-SRLwiki and QA-SRL 2.0 respectively. We see that the constrained models produce near-consistent gains over their unconstrained counterparts. Flan-T5 does not outperform the T5 model, but iterative prompting with inference significantly improves its performance. More importantly, without inference, the models produce highly inconsistent predictions with a lowest inconsistency percent of 34%! In addition, we also provide results on GPT-4 with the constraints verbalized as system instruction.111111GPT-4 results are meant to contextualize our results with a current SOTA system. The datasets we considered may have been utilized to train GPT-4. GPT-4 also has multiple orders of magnitude more parameters than the models we consider. For GPT-4, all the questions are provided at once. Refer Appendix D.1.2 for details.
@@
| System | Headq() | Heads() | () |
@@
| System | Headq() | Heads() | () |
@@ 6.1 Model Size
We study the performance and consistency of models over varying sizes. For SRL, we compare the small, base, large, 3 billion, and 11 billion T5 models. Figure 4 shows the Heads metric over T5 model sizes for QA-SRL 2.0, along with the inconsistency percent. For coreference, we compare the large, 3 billion and 11 billion variants of the Macaw model. We consider the All-Link inference for the constrained models. We report the F1 score and the inconsistency percent on the OntoNotes dataset in Figure 5. The inconsistency is shown only for the unconstrained models; the constrained ones are guaranteed to be consistent.
Across the datasets, we see that performance almost always increases and constraints provide steady gains. In QA-SRL 2.0, inconsistency reduces with increasing model size. The trend reverses for OntoNotes. Interestingly, with QA-SRL 2.0, a constrained model of a certain size often yields comparable, if not better, results as an unconstrained model of the immediately bigger model, showing that constraints might give performance gains equivalent to using a bigger model. Appendix C shows results for the other datasets.
@@ 6.2 Few-shot Prompting
So far, we have examined our framework in the zero-shot setting. Do the conclusions hold for few-shot settings as well? To answer this question, we use the Flan-T5-XL model which is known to yield promising results with even a few in-context examples.151515An interesting question: What counts as a shot? All questions in a structure or just one? We consider the latter. We conduct few-shot experiments for the SRL task on the QA-SRLwiki (Figure 6) and QA-SRL 2.0 (Figure 7). We observe a clear increase in performance with increasing shots for QA-SRLwiki. Recall that Flan T5, in a zero-shot setting for this dataset, underperforms compared to T5. Hence, the gains are not surprising: only the 5-shot constrained system beats the zero-shot constrained T5 model (38.75 Heads). The model also becomes more consistent with more examples. For QA-SRL 2.0, the performance and the consistency largely remain invariant with number of shots. Constrained systems produce more accurate structures than their unconstrained counterparts.
@@ Appendix A Detailed SRL Results
| System | Exactq() | Exacts() | Headq() | Heads() | () |
@@
| System | Exactq() | Exacts() | Headq() | Heads() | () |
@@ B.1 Zero Shot Coreference Results
The complete zero shot results for Macaw-3B and Flan T5-XL models are given in Table 8. One might notice that the Flan model is highly inconsistent as compared to the Macaw model. Note that in each case, the divisor is different since it is the number of times the antecedent is activated in equation 3. Examining the predictions, we saw that the value of the divisor for the Macaw models is roughly four times that of the Flan models. This implies that Flan-T5 tends to predict ‘No’ more often.
| Dataset | System | F1() | MUC() | B3() | CEAFe() | CoNLL() | () |
@@ B.2 Context Styles for Coreference Resolution
As mentioned in §5.2, for any given question pertaining to a mention-pair, we consider only the sentences where the constituent mentions occur. One can argue that surrounding context can play an important role in deciding whether or not two mentions are co-referent. Furthermore, intermediate context between the sentences can be pivotal in establishing otherwise loosely connected sentences appearing a few sentences apart. This begs the question about how much, if any, does providing the entire document help.161616For Ontonotes and GENIA, the document size went beyond memory which is why we do not test along this dimension. On a different note, mentions can be highlighted with asterisks(as used for WSC in Raffel et al. (2020)) to explicitly state where in the context they occur. This approach can help deal with pronominal mentions which are bound to repeat. We present results when the entire document (Full) acts as context, when mentions are highlighted (Hlght), and a combination of both (Full + Hlght) in comparison to the context style where only the relevant sentences are considered (Rel). We provide detailed results for performance and consistency over differing context styles in Table 9. We see that results are comparable across all context styles.
| Dataset | System | F1() | MUC() | B3() | CEAFe() | CoNLL() | () |
@@ Appendix C Additional Model Size Experiments
The analysis over model sizes for QA-SRLwiki is given in Figure 8. We see similar analysis as observed with QA-SRL 2.0 dataset. Performance improves with increasing model size while inconsistency reduces. Furthermore, constraints provide steady gains over their unconstrained counterparts.
@@ D.1.1 Iterative Prompting Template
In the iterative prompting template, we prompt the model one question at a time such that the answer for every preceding prompt is included in subsequent prompts. Assume a predicate in a sentence to have questions: {q1,q2,…,qn}. In the first step, we prompt the model for q1 using the prompt template mentioned in the previous sub-section. Next, we follow with the prompt the model for q2 such that the question and the generated answer for q1 are included in the prompt. Consequently, we use a different prompt template that contains this pair and an instruction that asks the answers to not overlap. Subsequently, the prompt for q3 will contain the question-answer pair q1 and q2, and so on. An example is shown in Figure 11. We use the order provided in the QA-SRL datasets for our experiments.
@@ Appendix F SRL Inference Worked-out Example
| Rank | Role | ||
| a | b | c | |
| 1 | “Elrond” (2) | “Aragorn” (5) | “Aragorn the sword” (5) |
| 2 | “Elrond gave” (1) | “Elrond” (3) | “the sword” (4) |